# Textured-GS: Gaussian Splatting with Spatially Defined Color and Opacity

### v1

@@ Abstract
In this paper, we introduce Textured-GS, an innovative method for rendering Gaussian splatting that incorporates spatially defined color and opacity variations using Spherical Harmonics (SH). This approach enables each Gaussian to exhibit a richer representation by accommodating varying colors and opacities across its surface, significantly enhancing rendering quality compared to traditional methods. To demonstrate the merits of our approach, we have adapted the Mini-Splatting architecture to integrate textured Gaussians without increasing the number of Gaussians. Our experiments across multiple real-world datasets show that Textured-GS consistently outperforms both the baseline Mini-Splatting and standard 3DGS in terms of visual fidelity. The results highlight the potential of Textured-GS to advance Gaussian-based rendering technologies, promising more efficient and high-quality scene reconstructions.
@@ 1 Introduction
Novel view synthesis is an active topic in Computer Vision and Graphics, focusing on generating accurate and realistic views from sparse input images with known camera parameters. Neural radiance fields (NeRF) [16] introduce a method to model a 3D scene by optimizing a continuous volumetric scene function through a fully connected deep network. This approach leverages implicit neural representations, extracting features like colors and geometry from a 3D scene by querying simple multi-layer perceptron networks with 5D inputs (spatial locations and camera views). The impressive realism achieved in free-viewpoint rendering by this learned mapping has sparked a surge of follow-up methods aimed at improving quality and speed, often incorporating regularization strategies. Notable examples include Mip-NeRF360 [1] and Zip-NeRF [2], which which excel in rendering quality, though they still require substantial training and rendering times. InstantNGP [18] enhances efficiency with a multiresolution hash encoding and a streamlined network architecture, reducing both training duration and computational demands during inference. Nonetheless, there are scenarios where NeRF’s rendering capability might not fully capture detailed scene elements, particularly in complex light interactions and reflections. Furthermore, NeRF-based methods often struggle to accurately represent empty spaces.
More recently, 3D Gaussian Splatting (3DGS) [11] has been introduced as a method for real-time rendering that significantly accelerates both the rendering process and scene optimization through a set of optimized 3D Gaussians, efficiently rasterizable on modern GPUs. Not only faster, its rendering quality matches or even surpasses that of leading NeRF implementations. Scenes are represented by millions of 3D Gaussians with specified position, rotation, scale, opacity, and color parameters, requiring significant storage and memory, rendering it impractical on devices with limited video memory like smartphones or head-mounted displays.
@@
Our objective is to enhance the textural representation of individual Gaussians to model color variations locally, even from a single viewing direction. By maintaining the existing SH framework and parameter allocation, we modify the parameterization scheme. This adjustment allows each 3D Gaussian to display different colors under various viewing angles and across different areas of the Gaussian ellipsoidal surface when viewed from the same angle. Additionally, we integrate an opacity channel into the SH framework, enabling the modeling of opacity variations along the Gaussian surface, thereby enriching visual complexity and increasing the realism of rendered scenes.
@@
  • We extend this technique to model opacity variations on Gaussian ellipsoidal surfaces, allowing for deviations from the standard ellipsoidal shape. This includes creating “uni-directional Gaussians” that can be opaque on one side and transparent on the other.
  • We validated our approach by applying it to 3D Gaussian scenes pre-optimized using Mini-splatting [5]. Our method achieves state-of-the-art rendering quality on several established datasets while requiring significantly fewer Gaussians compared to standard 3D Gaussian Splatting (3DGS) [11].
@@ 2.1 Novel View Synthesis
Introduced by Mildenhall et al. [16], Neural Radiance Fields (NeRF) leverage a fully-connected neural network to model a volumetric scene function continuously, enabling precise capture of color and density throughout 3D space. NeRF generates new perspectives with notable detail and realism by processing these predictions through a differentiable rendering framework along the viewer’s line of sight. Despite its impressive output quality, NeRF’s main drawback is its computational intensity and slow rendering times, which have prompted numerous subsequent studies aimed at overcoming these challenges [7, 3, 18, 6, 15, 22, 13].
Key advancements in this area include Instant-NGP [18], which employs a neural hash grid to markedly decrease computational load while sustaining output quality. Plenoxels [6] utilize a sparse, learnable grid that eschews traditional neural networks, significantly expediting both training and inference. Mip-NeRF 360 [1] focuses on improving renderings within 360-degree environments, tackling the challenges associated with view variability and inconsistent lighting. Furthermore, Zip-NeRF [2] innovates on network compression methods to enhance rendering fidelity without sacrificing the quality of the produced images.
@@
Developed by Kerbl et al. as an efficient alternative to traditional volume rendering technuques, 3DGS [11] represents scene through a sparse set of 3D Gaussians. Each Gaussian in this representation is characterized by its position, color, and covariance matrix. 3DGS stands out due to its differentiable rendering process, allowing it to be optimized directly from photometric observations. This technique not only speeds up the rendering process but also enhances the adaptability of the representation, making it particularly effective for real-time applications such as VR and AR where rapid rendering is crucial. Although the 3D Gaussian Splatting achieves real-time rendering, there is improvement space in terms of lower computational requirements and better point distribution [23].
Various 3DGS compression techniques employ vector quantization to cluster multi-dimensional data into a finite set of representations [14, 19, 20, 8, 4]. Specifically, Niedermayr et al. [20] utilizes vector clustering to compactly encode color and geometric attributes into two codebooks to reduce redundancy. Similarly, the EAGLES [8] applies quantization across all attributes of each Gaussian and show that the quantization of opacity leads to fewer floaters or visual artifacts. However, these techniques often do not address the suboptimal distribution of Gaussians, which tends to result in local minima following compression.
Rather than simply compressing existing 3D Gaussian Splatting systems, several initiatives aim to improve Gaussian distribution [5, 4, 17, 10]. LightGaussian [4] reduces the number of Gaussians by pruning those with lower importance scores and employs an octree-based method for compressing positions. Jo et al. [10] developed a strategy to both compress 3DGS and enhance computational efficiency by eliminating non-essential Gaussians. Meanwhile, Mini-Splatting [5] enhances the rendering process by utilizing more effective Gaussian splats. This method samples Gaussians based on importance score instead of pruning them to avoid artifacts.
@@
Mini-Splatting [5] 0.822 27.34 0.217 0.49 0.849 23.18 0.202 0.20 0.908 29.98 0.253 0.35
@@
Table 1: Quantitative evaluation of our Spatial-GS and previous works in three commonly-used metrics: SSIM, PSNR and LPIPS. Num represent number of Gaussians (million).
@@ 5.1 Datasets
We conducted experiments across three real-world datasets: Mip-NeRF360 [1], Tanks and Temples [12], and Deep Blending [9]. To ensure consistency and fairness in our evaluations, we adhered to the same processing methodologies used in the 3DGS and Mini-Splatting studies, including identical scene selection, train/test splits, and image resolution settings. Specifically, for each dataset, we implemented a train/test division following the recommendations of Mip-NeRF 360 [1], selecting every eighth photo for testing to facilitate consistent and meaningful comparisons. Our evaluations utilized standard error metrics widely used in the field, such as PSNR, LPIPS, and SSIM.
@@ 5.2 Implementation Details
We implemented Textured-GS using PyTorch and integrate it into the optimization pipeline of 3D Gaussian Splatting (3DGS) [11]. As discussed previously, we removed the module for adaptive control of Gaussians. We also modified the Gaussian rasterization module within the 3DGS framework to facilitate the rendering of textured Gaussians and update SH coefficients through gradient descent.
Our experiments were conducted on an Ubuntu 20.04 platform with an NVIDIA RTX 4090 GPU. Starting with the output Gaussians from Mini-Splatting [5], we trained our model for 14,000 iterations, applying learning rate of 0.0025 for color texture and 0.005 for opacity texture, respectively. It is important to note that view dependent color was not loaded as part of our input due to a fundamentally different Gaussian setup. We opted for the square root of the scale parameters, $\sqrt{s}$, for the ellipsoid in Equation 8, because larger ellipsoids tended to yield better performance, though this does not alter the actual Gaussian size in splatting. All other parameters were aligned with those used in Mini-Splatting and 3DGS to ensure a fair comparison.
Method Num FPS Train Time Memory
3DGS [11] 6.03M 92 26m55s 11.77GB
Mini.-D [5] 6.02M 105 23m45s 11.57GB
Mini. [5] 0.53M 416 11m33s 5.46GB
Ours 0.53M 52 21m30s 4.53GB
@@
Table 3: Ablation study assessing the impact of varying SH degrees on color and opacity using the Mip-NeRF 360 dataset. Employing SH degree zero for both components reduces the baseline to Mini-Splatting with mono-color Gaussians [5]. The highest SH degrees yield the best performance, clearly demonstrating the benefits of incorporating color and opacity textures.
@@ 5.3 Results
Rendering Quality Table 1 presents a quantitative evaluation across three different real-world dataset. We compare our Textured-GS with the baseline method Mini-Splatting, 3DGS, and other related algorithms. In most categories, our Textured-GS outperforms both 3DGS and Mini-Splatting. The NeRF-based method Zip-NeRF [2], currently the state-of-the-art on Mip-NeRF 360 dataset, shows superior performance. Mini-Splatting-D [5], a denser version of Mini-Splatting, surpasses ours method in some categories but uses a greater number of Gaussians than 3DGS, contradicting our goal of efficiency.
@@
Figure 4 further plots the performance of our approach compared to Mini-Splatting [5] across various numbers of Gaussians. Our method consistently outperforms Mini-Splatting at each scale. Models with a fewer number of Gaussians are generated by using a lower sampling rate within the Mini-Splatting framework.
@@
Resource Consumption We assessed the resource consumption of various rendering methods, focusing on the number of Gaussians, frames per second (FPS), training time, and maximum allocated memory during training. Table 2 summarizes the comparative performance and efficiency of these methods. The results show a clear reduction in memory usage. It is worth noting that both the rendering speed and training time have significant room for optimization.
@@ 5.4 Ablation Study
We conducted an ablation study to evaluate the impact of two key components of our proposed method: textured color and textured opacity. For each component, we gradually increased the Spherical Harmonics (SH) degree from zero to three. Table 3 summarizes the results, tested on the Mip-NeRF 360 dataset. Our method, employing the third degree level for both color and opacity, achieves the best results, demonstrating incremental improvements as the SH degree level increases.
@@ 5.5 Limitations
Our method enhances the capabilities of each Gaussian in the Mini-Splatting framework [5] by introducing textured color and opacity, leading to improved rendering quality. However, this approach has limitations, notably the increase in computational cost due to additional calculations for ray-ellipsoid intersections, which result in longer training and rendering times. We believe this limitation can be addressed by optimizing the rendering process and integrating the sampling for color and opacity into a single step, thereby streamlining computations and reducing overhead.
To isolate the impact of optimizing color and opacity textures, we maintained the locations, sizes, and orientations of Gaussians as generated by Mini-Splatting unchanged. Additionally, we removed the module for adaptive control of Gaussians, as found in 3DGS. Optimizing all these parameters in an end-to-end manner is likely to enhance the performance of our algorithm. As part of our future work, we plan to develop a comprehensive framework that starts with a structure-from-motion (SFM) point cloud, employs mechanisms for both densifying and simplifying Gaussians, and ultimately produces a set of textured Gaussians that optimally represent the scenes.
@@ 6 Conclusion
In conclusion, our study introduces Textured-GS, an innovative rendering method that leverages spherical harmonics to introduce spatially defined color and opacity variations within each Gaussian. This approach significantly enhances the visual quality of renderings, providing a more detailed and flexible solution than traditional Gaussian splatting methods. To validate the effectiveness of our method, we integrated Textured-GS within the Mini-Splatting framework without increasing the number of Gaussians used. Our comprehensive evaluation across three real-world datasets demonstrates that Textured-GS consistently outperforms both the baseline Mini-Splatting and standard 3DGS in rendering quality. We also observed that our method efficiently addresses complex scenes that typically challenge Gaussian-based approaches, such as sharp edges and detailed structures. Moving forward, we plan to develop a fully optimized end-to-end framework that incorporates adaptive control of Gaussians, which we anticipate will further enhance the capabilities and efficiency of our approach, making it even more suitable for high-fidelity rendering tasks.

### v2

@@ Abstract
In this paper, we introduce Textured-GS, an innovative method for rendering Gaussian splatting that incorporates spatially defined color and opacity variations using Spherical Harmonics (SH). This approach enables each Gaussian to exhibit a richer representation by accommodating varying colors and opacities across its surface, significantly enhancing rendering quality compared to traditional methods. To demonstrate the merits of our approach, we adapted the Mini-Splatting architecture to integrate textured Gaussians without increasing the number of Gaussians. Our experiments across multiple real-world datasets show that Textured-GS consistently outperforms both the baseline Mini-Splatting and standard 3DGS in terms of visual fidelity. These results highlight the potential of Textured-GS to advance Gaussian-based rendering technologies, promising more efficient and high-quality scene reconstructions.
@@ 1 Introduction
Novel view synthesis is an active topic in Computer Vision and Graphics, focusing on generating accurate and realistic views from sparse input images with known camera parameters. Neural radiance fields (NeRF) [17] introduce a method to model a 3D scene by optimizing a continuous volumetric scene function through a fully connected deep network. This approach leverages implicit neural representations, extracting features like colors and geometry from a 3D scene by querying simple multi-layer perceptron networks with 5D inputs (spatial locations and camera views). The impressive realism achieved in free-viewpoint rendering by this learned mapping has sparked a surge of follow-up methods aimed at improving quality and speed, often incorporating regularization strategies. Notable examples include Mip-NeRF360 [1] and Zip-NeRF [2], which excel in rendering quality, although they still require substantial training and rendering times. InstantNGP [19] enhances efficiency with multiresolution hash encoding and a streamlined network architecture, reducing both the training duration and the computational demands during inference. However, there are scenarios where NeRF’s rendering capability might not fully capture detailed scene elements, particularly in complex light interactions and reflections. Furthermore, NeRF-based methods often struggle to accurately represent empty spaces.
More recently, 3D Gaussian Splatting (3DGS) [11] has been introduced as a method for real-time rendering that significantly accelerates both the rendering process and scene optimization through a set of optimized 3D Gaussians, efficiently rasterizable on modern GPUs. In addition to its speed, 3DGS achieves a rendering quality that matches or even surpasses the leading NeRF implementations. Scenes are represented by millions of 3D Gaussians with specified position, rotation, scale, opacity, and color parameters, requiring significant storage and memory, rendering it impractical on devices with limited video memory like smartphones or head-mounted displays.
@@
Our objective is to enhance the textural representation of individual Gaussians to model color variations locally, even from a single viewing direction. By maintaining the existing SH framework and parameter allocation, we modify the parameterization scheme. This adjustment allows each 3D Gaussian to display different colors at various viewing angles and across different areas of the Gaussian ellipsoidal surface when viewed from the same angle. Additionally, we integrate an opacity channel into the SH framework, enabling the modeling of opacity variations along the Gaussian surface, thereby enriching visual complexity and increasing the realism of rendered scenes.
@@
  • We extended this technique to model opacity variations on Gaussian ellipsoidal surfaces, allowing for deviations from the standard ellipsoidal shape. This includes creating “unidirectional Gaussians” that can be opaque on one side and transparent on the other.
  • We validated our approach by applying it to pre-optimized 3D Gaussian scenes using mini-splatting [5]. Our method achieves comparable rendering quality on established datasets while requiring significantly fewer number of Gaussians compared to standard 3D Gaussian Splatting (3DGS) [11].
@@ 2.1 Novel View Synthesis
Introduced by Mildenhall et al. [17], Neural Radiance Fields (NeRF) leverage a fully connected neural network to continuously model a volumetric scene function, enabling precise capture of color and density throughout 3D space. NeRF generates new perspectives with notable detail and realism by processing these predictions through a differentiable rendering framework along the viewer’s line of sight. Despite its impressive output quality, the main drawback of NeRF is its computational intensity and slow rendering times, which have led to numerous subsequent studies aimed at overcoming these challenges [7, 3, 19, 6, 16, 23, 14].
Key advancements in this area include Instant-NGP [19], which employs a neural hash grid to markedly decrease computational load while sustaining output quality. Plenoxels [6] utilize a sparse, learnable grid that eschews traditional neural networks, significantly expediting both training and inference. Mip-NeRF 360 [1] focuses on improving renderings within 360-degree environments, tackling the challenges associated with view variability and inconsistent lighting. Furthermore, Zip-NeRF [2] innovates in network compression methods to enhance rendering fidelity without sacrificing the quality of the produced images.
@@
Developed by Kerbl et al. as an efficient alternative to traditional volume rendering techniques, 3DGS [11] represents scene through a sparse set of 3D Gaussians. Each Gaussian in this representation is characterized by its position, color, and covariance matrix. 3DGS stands out due to its differentiable rendering process, allowing it to be optimized directly from photometric observations. This technique not only speeds up the rendering process but also enhances the adaptability of the representation, making it particularly effective for real-time applications such as VR and AR where rapid rendering is crucial. Although the 3D Gaussian Splatting achieves real-time rendering, there is improvement space in terms of lower computational requirements and better point distribution [24].
Various 3DGS compression techniques employ vector quantization to cluster multi-dimensional data into a finite set of representations [15, 20, 21, 8, 4]. Specifically, Niedermayr et al. [21] utilizes vector clustering to compactly encode color and geometric attributes into two codebooks to reduce redundancy. Similarly, the EAGLES [8] applies quantization across all attributes of each Gaussian and show that the quantization of opacity leads to fewer floaters or visual artifacts. However, these techniques often do not address the suboptimal distribution of Gaussians, which tends to result in local minima following compression. We believe that similar compression techniques can also be applied on Textured-GS in the future work, as there are four channels of SH coefficients to be compressed.
Rather than simply compressing existing 3D Gaussian Splatting systems, several initiatives aim to improve Gaussian distribution [5, 4, 18, 10]. LightGaussian [4] reduces the number of Gaussians by pruning those with lower importance scores and employs an octree-based method for compressing positions. Jo et al. [10] developed a strategy to both compress 3DGS and enhance computational efficiency by eliminating non-essential Gaussians. Meanwhile, Mini-Splatting [5] enhances the rendering process by utilizing more effective Gaussian splats. This method samples Gaussians based on importance score instead of pruning them to avoid artifacts. It reduced the scenario where multiple Gaussians are stacked to form one color, which aligns with our objective of applying textures to individual Gaussians. Therefore, it has been selected as our baseline method.
@@
Mini-Splatting (30K) [5] 0.822 27.34 0.217 0.49 0.835 23.18 0.202 0.20 0.908 29.98 0.253 0.35
Mini-Splatting (44K) [5] 0.822 27.34 0.213 0.49 0.840 23.34 0.203 0.20 0.906 29.80 0.250 0.35
@@
Table 1: Quantitative evaluation of our Textured-GS and previous works in three commonly-used metrics: SSIM, PSNR and LPIPS. Num represent number of Gaussians (million).
@@ 5.1 Datasets
We conducted experiments across three real-world datasets: Mip-NeRF360 [1], Tanks and Temples [13], and Deep Blending [9]. To ensure consistency and fairness in our evaluations, we adhered to the same processing methodologies used in the 3DGS and Mini-Splatting studies, including identical scene selection, train/test splits, and image resolution settings. Specifically, for each dataset, we implemented a train/test division following the recommendations of Mip-NeRF 360 [1], selecting every eighth photo for testing to facilitate consistent and meaningful comparisons. Our evaluations used standard error metrics widely used in the field, such as PSNR, LPIPS, and SSIM.
@@ 5.2 Implementation Details
We implemented Textured-GS using PyTorch and integrate it into the 3D Gaussian splatting (3DGS) [11] optimization pipeline. As discussed previously, we removed the module for adaptive control of Gaussians. We also modified the Gaussian rasterization module within the 3DGS framework to facilitate the rendering of textured Gaussians and update SH coefficients through gradient descent.
Our experiments were carried out on an Ubuntu 20.04 platform with an NVIDIA RTX 4090 GPU. Starting with the output Gaussians from Mini-Splatting [5], we trained our model for 14,000 iterations, applying a learning rate of 0.0025 for the color texture and 0.005 for the opacity texture, respectively. It is important to note that view dependent color was not loaded as part of our input due to a fundamentally different Gaussian setup. We opted for the square root of the scale parameters, $\sqrt{s}$, for the ellipsoid in Equation 8, because larger ellipsoids tend to yield better performance, though this does not alter the actual Gaussian size in splatting. All other parameters were aligned with those used in Mini-Splatting and 3DGS to ensure a fair comparison.
Method Num FPS Train Time Memory File Size
3DGS [11] 6.03M 92 26m55s 11.77GB 1.5GB
Mip-Splatting [25] 7.88M 62 40m43s 15.21GB 2.0GB
Mini.-D [5] 6.02M 105 23m45s 11.57GB 1.5GB
Mini. [5] 0.53M 416 11m33s 5.46GB 132MB
Ours 0.53M 161 10m52s 4.53GB 164MB
@@
Table 3: Ablation study assessing the impact of varying SH degrees on color and opacity using the Mip-NeRF 360 dataset. Employing SH degree zero for both components reduces the baseline to Mini-Splatting with monocolored Gaussians [5]. The highest SH degrees yield the best performance, clearly demonstrating the benefits of incorporating color and opacity textures.
@@ 5.3 Results
Rendering Quality Table 1 presents a quantitative evaluation across three different real-world dataset. We compare our Textured-GS with the baseline method Mini-Splatting, 3DGS, and other related algorithms. In most categories, our Textured-GS outperforms both 3DGS and Mini-Splatting. The NeRF-based method Zip-NeRF [2], currently the state-of-the-art on the Mip-NeRF 360 dataset, shows superior performance. Mini-Splatting-D [5], a denser version of Mini-Splatting, and Mip-Splatting [25], surpass our method in some categories but both use greater numbers of Gaussians than 3DGS, contradicting our goal of efficiency.
@@
Figure 4 further plots the performance of our approach compared to Mini-Splatting [5] on various numbers of Gaussians. Our method consistently outperforms the Mini-Splatting at each scale. Models with a fewer number of Gaussians are generated by using a lower sampling rate within the Mini-Splatting framework.
@@
Resource Consumption We assessed the resource consumption of various rendering methods, focusing on the number of Gaussians, frames per second (FPS), training time, and the maximum memory allocated during training. Table 2 summarizes the comparative performance and efficiency of these methods. The results show that our approach uses less training memory and slightly more storage size than the baseline method with the same number of Gaussians. The reason is that no depth calculation is required in our framework, but four channels of SH coefficients are stored instead of three. It should be noted that both the rendering speed and the training time still have room for optimization.
@@ 5.4 Ablation Study
We conducted an ablation study to evaluate the impact of two key components of our proposed method: textured color and textured opacity. For each component, we gradually increased the degree of Spherical Harmonics (SH) from zero to three. Table 3 summarizes the results, tested on the Mip-NeRF 360 dataset. Our method, employing the third degree level for both color and opacity, achieves the best results, demonstrating incremental improvements as the SH degree level increases.
@@ 5.5 Limitations
Our method enhances the capabilities of each Gaussian in the Mini-Splatting framework [5] by introducing textured color and opacity, leading to improved rendering quality. However, this approach has limitations, notably the increased computational cost due to additional calculations for ray-ellipsoid intersections, resulting in longer training and rendering times. We believe this limitation can be addressed by optimizing the rendering process and integrating the sampling for color and opacity into a single step, thereby streamlining computations and reducing overhead.
To isolate the impact of optimizing color and opacity textures, we kept the locations, sizes, and orientations of Gaussians as generated by Mini-Splatting unchanged. Additionally, we removed the adaptive control module for Gaussians, as found in 3DGS. Optimizing all these parameters in an end-to-end manner is likely to enhance the performance of our algorithm. As part of our future work, we plan to develop a comprehensive framework that starts with a structure-from-motion (SFM) point cloud, incorporates mechanisms for both densifying and simplifying Gaussians, and ultimately produces a set of textured Gaussians that optimally represent the scenes.
@@ 6 Conclusion
In conclusion, our study introduces Textured-GS, an innovative rendering method that leverages spherical harmonics to introduce spatially defined color and opacity variations within each Gaussian. This approach significantly enhances the visual quality of renderings, offering a more detailed and flexible solution compared to traditional Gaussian splatting methods. To validate the effectiveness of our method, we integrated Textured-GS into the Mini-Splatting framework without increasing the number of Gaussians. Our comprehensive evaluation across three real-world datasets demonstrates that Textured-GS consistently outperforms both the baseline Mini-Splatting and standard 3DGS in rendering quality. Additionally, we observed that our method efficiently addresses complex scenes that typically challenge Gaussian-based approaches, such as sharp edges and detailed structures. Moving forward, we plan to develop a fully optimized end-to-end framework that incorporates adaptive control of Gaussians, which we anticipate will further enhance the capabilities and efficiency of our approach, making it even more suitable for high-fidelity rendering tasks.